OpenStudy (anonymous):

What is the y-intercept of the function y = 0.5(8)x? 8 0.5 1

OpenStudy (anonymous):

y=0.5(8)^x for y intercept x=0 so, y=0.5(8)^0=0.5*1=0.5 so, 0.5 is the y intercept